Huana spent $28 on writing on writing utensils, pens cost $4 dollars and $2 pencils. She bought a total of 8. How many did she buy?

Found 2 solutions by JulietG, Alan3354:
Answer by JulietG(1812)

This can be done a few ways. Since it's a small number, we'll do it the easiest way - estimation.
What is the maximum number of pens that can be bought for $28? 28/4 = 7
That's not enough items. What if we buy 6? 6*4 = 24. That leaves us 2 items short. 2*2 = 4. 24+4=28.
6 pens and 2 pencils

Answer by Alan3354(69443)

Huana spent $28 on writing on writing utensils, pens cost $4 dollars and $2 pencils. She bought a total of 8. How many did she buy?
-------------
P = # of pens
C = # of pencils
---
P + C = 8
4P + 2C = 28